OpenStudy (anonymous):

integral of 4^(3x+1)

hartnn (hartnn):

try u= 3x+1 du =...

OpenStudy (anonymous):

\[\int\limits_{}^{}4^{3x+1}dx\]

OpenStudy (anonymous):

i actually did ! but for some reason the 4 is throwing me off.. \[\frac{ 1 }{ 3 }\int\limits_{}^{}4^{u}du\]

hartnn (hartnn):

integral of a^x is ... ? there's a standard formula

OpenStudy (anonymous):

\[\frac{ a ^{x} }{ \ln a }+ C\]

hartnn (hartnn):

yes here a = 4

OpenStudy (anonymous):

mkay so my answer that i got is….\[\frac{ 4^{3x+1} }{ 3\ln 4 } + C\]

hartnn (hartnn):

thats absolutely correct! :)

OpenStudy (anonymous):

:O Yes! :)